OpenStudy (anonymous):

To find the distance from the house at A to the house at B, a surveyor measures the angle ACB, which is found to be 70°, and then walks off the distance to each house, 50 feet(A to C) and 70 feet(C to B), respectively. How far apart are the houses? (answer to 1 decimal place) in feet

OpenStudy (amistre64):

so we got a ABC tri AC =50 ;BC = 70 and angle c=70 degrees law of cosines

OpenStudy (amistre64):

(AB)^2 = 50^2 + 70^2 -2(50)(70) cos(70)

OpenStudy (amistre64):

google gives me: sqrt((50^2) + (70^2) - (2 * 50 * 70 * cos(70))) = 54.4680234

OpenStudy (amistre64):

54.5

OpenStudy (anonymous):

the answer is wrong :)

OpenStudy (anonymous):

would the answer be 85.8?

OpenStudy (amistre64):

i got the 50 and 70 backwards in the diagram; but that doesnt affect the answer

OpenStudy (anonymous):

ye true. when i calculated it i got 85.8? calc error?

OpenStudy (amistre64):

2500 + 4900 - 7000(cos(70)) = 54.4680

OpenStudy (amistre64):

well; the square root of that equals 54.4680

OpenStudy (amistre64):

54.4; try that?

OpenStudy (anonymous):

sry im gettin 70.75

OpenStudy (amistre64):

then your hitting the wrong button lol

OpenStudy (amistre64):

did you typo the original question you posted?

OpenStudy (anonymous):

nah i typed it correctly. could it be the picture u drew?

OpenStudy (amistre64):

the picture? lol.... that doesnt affect the math

OpenStudy (amistre64):

i labeld it and everything lol

OpenStudy (anonymous):

i put in 70.8. that is the correct answer

OpenStudy (amistre64):

what formula did you use?

OpenStudy (anonymous):

the exact same that you used. i think its just a calc error

OpenStudy (amistre64):

i used afew different calculators :)

OpenStudy (anonymous):

haha lol. hey do u know how i can put a pic with my question?

OpenStudy (amistre64):

if you have it in a file on your comluter; just use the 'attach file' button located at the bottom of this reply window

OpenStudy (amistre64):

wolfram says your right lol

OpenStudy (anonymous):

lol but i give u all the credit hahaha

OpenStudy (amistre64):

my calcs were in radians disguised as degrees lol

OpenStudy (amistre64):

70 rads is not the same as 70 degrees

OpenStudy (anonymous):

hahah. i always do that mistake

OpenStudy (amistre64):

there should be a "are you sure your not forgetting something" option that pops up lol

OpenStudy (anonymous):

haha if only!! that would be epic.

OpenStudy (anonymous):

mate which country r u from?

OpenStudy (amistre64):

usa... down here in florida

OpenStudy (anonymous):

wow thats a long way from here!! im down under at australia

OpenStudy (amistre64):

yeah; but it shorter if you go thru hyperspace ;)

OpenStudy (anonymous):

hahahaha true that!!! you better invent that one day